# PLUTO-4: Frontier Pathology Foundation Models

### v1

@@ 1 Introduction
The training methodology builds upon the original PLUTO workflow (Juyal et al., 2024), maintaining a self-supervised pretraining framework derived from DINOv2 (Oquab et al., 2024) with multi-resolution image sampling. PLUTO-4 extends this framework with improved distributed training stability, higher-precision optimization, and extended compute scaling, enabling consistent convergence for both compact and frontier-scale architectures.
@@ 3.1 Training Architecture Design
The compact encoder PLUTO-4S is designed to support multiple levels of representational granularity across diverse pathology tasks. Different problems, such as cell segmentation versus slide-level classification, require varying receptive field sizes and granular context to capture relevant morphologic context. To accommodate this, we adopt the FlexiViT setup (Beyer et al., 2023) similar to PLUTO (Juyal et al., 2024), allowing the same backbone to operate on variable patch-token sizes. During pretraining, patch-token sizes from [8, 16, 32] are sampled dynamically across training iterations, exposing the model to multi-scale input structure without separate backbone training. Unlike the absolute positional embeddings used in PLUTO (Juyal et al., 2024), PLUTO-4 employs two-dimensional rotary positional embeddings (2D-RoPE) (Su et al., 2021), which provide relative positional encoding in both spatial dimensions and improve stability at large sequence lengths. This modification is consistent with the positional encoding formulation used in (Siméoni et al., 2025) and other large-scale vision transformers.
@@ 4 Results and Evaluation
CRC (Test) 96.8 96.2 96.7 97.1
PANDA-Small 66.6 68.0 66.4 70.5
@@ 4.3.4 Slide-Level Prediction
On the PANDA Gleason grading benchmark, PLUTO-4G demonstrates strong performance, surpassing models such as Virchow-2, UNI2-H, and Prov-Gigapath, performing on par with H-Optimus-0, and slightly below Atlas.(Table 1).
@@
Table 3: Performance results for PLUTO-4: CLS vs. CLS + Mean(Patch-Token) across patch sizes and models. ViT-S models trained with FlexiViT achieve strong performance while allowing configuration of the best patch-token size [8 / 16] for the task. Using CLS + Patch embeddings yields modest improvements across most classification tasks.
@@
Benchmark datasets are setup and evaluated using EVA, with the exception of HEST, Derm-2K (proprietary). For segmentation datasets, CLS + Mean(Patch-Token) features are not applicable, as segmentation decoders utilize the entire set of Patch-Token features rather than pooled representations.
@@ 4.3.5 CLS + Mean(Patch-Token) Concatenation
Across eight benchmark datasets, PLUTO-4G achieves or matches state-of-the-art performance on nearly every task, outperforming or equaling other external models such as Atlas, Virchow-2, and H-Optimus-0. Specifically, PLUTO-4G attains the highest correlation on HEST (43.2) and top accuracy on MHIST (87.9), BACH (93.2), and PCAM (95.2). For the slide-level Gleason grading Panda-small, PLUTO-4G achieves comparable performance to other models, while Atlas retains the best performance like it has with CLS. Overall, these results demonstrate that PLUTO-4G’s representations remain highly competitive even when evaluated under alternative embedding strategies used by other pathology foundation models.
@@ 4.4 Comparison of PLUTO Models
For PLUTO-4G, we observe moderate gains on several classification tasks, including HEST (+0.5 $r$), MHIST (+0.4 pp), PCAM, and CRC (+0.4 pp), while datasets such as BACH and Gleason show marginal declines. Across model variants, the improvements are not uniform but indicate a consistent trend: performance increases on 7 of 8 datasets for PLUTO-4S-8, 5 of 8 for PLUTO-4S-16, and 5 of 8 for PLUTO-4G. These results suggest that incorporating mean patch-token information can enhance discriminative ability for certain tasks, particularly those requiring finer-grained morphological context.

### v2

@@ 1 Introduction
The training methodology builds upon the original PLUTO workflow (Juyal et al., 2024), maintaining a self-supervised pretraining framework derived from DINOv2 (Oquab et al., 2024) with multi-resolution image sampling. PLUTO-4 extends this framework with improved distributed training stability, higher-precision optimization, architectural improvements and extended compute scaling, enabling consistent convergence for both compact and frontier-scale architectures.
@@ 3.1 Training Architecture Design
The compact encoder PLUTO-4S is designed to support multiple levels of representational granularity across diverse pathology tasks. Different problems, such as cell segmentation versus slide-level classification, require varying receptive field sizes and granular context to capture relevant morphologic context. To accommodate this, we adopt the FlexiViT setup (Beyer et al., 2023) similar to PLUTO (Juyal et al., 2024), allowing the same backbone to operate on variable patch-token sizes. During pretraining, patch-token sizes from [8, 16, 32] are sampled dynamically across training iterations, exposing the model to multi-scale input structure without separate backbone training. Unlike the absolute positional embeddings used in PLUTO (Juyal et al., 2024), PLUTO-4S employs two-dimensional rotary positional embeddings (2D-RoPE) (Su et al., 2021), which provide relative positional encoding in both spatial dimensions and improve stability at large sequence lengths. This modification is consistent with the positional encoding formulation used in (Siméoni et al., 2025) and other large-scale vision transformers.
@@ 4 Results and Evaluation
CRC 96.8 96.2 96.7 97.1
PANDA-Small (Test) 66.6 68.0 66.4 70.5
@@ 4.3.4 Slide-Level Prediction
On the PANDA-Small Gleason grading benchmark, PLUTO-4G demonstrates strong performance, surpassing models such as Virchow-2, UNI2-H, and Prov-Gigapath, performing on par with H-Optimus-0, and slightly below Atlas.(Table 1).
@@
Table 3: Performance results for PLUTO-4: CLS vs. CLS + Mean(Patch-Token) across patch sizes and models. ViT-S models trained with FlexiViT achieve strong performance while allowing configuration of the best patch-token size [8 / 16] for the task. Patch-Token size 8 significantly improves 4S performance on tasks needing fine-grained context.
Using CLS + Patch embeddings yields modest improvements across most classification tasks.
@@
Benchmark datasets are setup and evaluated using EVA, with the exception of HEST. For segmentation datasets, CLS + Mean(Patch-Token) features are not applicable, as segmentation decoders utilize the entire set of Patch-Token features rather than pooled representations.
@@ 4.3.5 CLS + Mean(Patch-Token) Concatenation
Across eight benchmark datasets, PLUTO-4G achieves or matches state-of-the-art performance on nearly every task, outperforming or equaling other external models such as Atlas, Virchow-2, and H-Optimus-0. Specifically, PLUTO-4G attains the highest correlation on HEST (0.432) and top accuracy on MHIST (87.9), BACH (93.2), and PCAM (95.2). For the slide-level Gleason grading PANDA-Small, PLUTO-4G achieves comparable performance to other models, and slightly below Atlas. Overall, these results demonstrate that PLUTO-4G’s representations remain highly competitive even when evaluated under alternative embedding strategies used by other pathology foundation models.
@@ 4.4 Comparison of PLUTO Models
For PLUTO-4G, we observe moderate gains on several classification tasks, including HEST (+0.005 $r$), MHIST (+0.4 pp), PCAM, and CRC (+0.4 pp), while datasets such as BACH and Gleason show marginal declines. Across model variants, the improvements are not uniform but indicate a consistent trend: performance increases on 7 of 8 datasets for PLUTO-4S-8, 5 of 8 for PLUTO-4S-16, and 5 of 8 for PLUTO-4G. These results suggest that incorporating mean patch-token information can enhance discriminative ability for certain tasks, particularly those requiring finer-grained morphological context.